OpenStudy (anonymous):

The sum of two consecutive integers is at most 223. What is the larger of the two integers?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

The answer is 112. Consecutive integers are numbers that come after the previous number. 111 + 112 = 223.

OpenStudy (anonymous):

A figure skater needs a total score of at least 90 to move on to the next round. Her total score is the average of four judges’ scores. The first three judges’ scores were 83, 88, and 92. The figure skater found out she made it to the next round. What is the minimum score the fourth judge could have given her?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Easy. Find the mean of the three numbers. 83 + 88 + 92 = 263 Divide by 3: 263 / 3 = 87.67 You need to find a number that will get her to 90. Another way to ask this is: x divided by 4 = 90. To solve multiply 4 to both sides: 4 * 90 = 360 So 360. So we need to find a fourth number that will add up to 60. The sum of the 3 numbers is 263. 360 - 263 = 97. So the judge had to give her a 97 at the least.

OpenStudy (anonymous):

Are you in K12 @Xgloves11 ?

OpenStudy (anonymous):

@CrazyCountryGirl Yup